# Reliable Grid Forecasting: State Space Models for Safety-Critical Energy Systems

### v5

@@ Appendix A Weather Covariates
Table 5 lists the 8 meteorological covariates used in weather-integrated models. All weather data are sourced from NOAA Integrated Surface Database (ISD) stations within each utility service territory, aggregated to hourly resolution.
@@ Appendix B Hyperparameter Configuration
Table 6 provides complete hyperparameter specifications for all models.
| Parameter | S-Mamba | PowerMamba | Mamba-ProbTSF | iTransformer |
| Embedding dim () | 128 | 128 | 128 | 512 |
| State dim () | 16 | 16 | 16 | — |
| Conv kernel () | 4 | 4 | 4 | — |
| Expansion factor | 2 | 2 | 2 | — |
| Encoder layers | 2 | 2 | 2 | 10 |
| Attention heads | — | — | — | 8 |
| Dropout | 0.1 | 0.1 | 0.1 | 0.1 |
| Bidirectional | Yes | Yes | Yes | — |
| Total parameters | 16.4M | 2.5M | 16.4M | 50.0M |
@@ Appendix C Multi-Seed Robustness
To assess robustness to random initialization, we trained PowerMamba + Weather with seeds . The 24-hour MAPE on CA ISO-TAC was (mean std), confirming that reported results are stable across initializations. Full multi-seed results for all architectures will be provided in the final version.
@@ Appendix D Per-Utility Results
Table 7 presents 24-hour MAPE for each TAC area, demonstrating consistent performance across diverse grid regions.
@@
Smaller utilities (SDGE-TAC, TIDC) exhibit higher MAPE due to increased load volatility relative to system size. Weather integration tends to improve accuracy, with the largest relative gains in smaller, more volatile territories; however, improvements are not uniform across all regions (Table 7).
@@ Appendix E Sensitivity to Weather Forecast Errors
Table 8 demonstrates that the model retains its performance advantage even with moderate forecast errors. At (a typical error range for day-ahead weather forecasts), the performance degrades by less than 10%, remaining competitive with the non-weather-integrated baselines. This confirms that the benefits of weather integration largely persist under operational weather uncertainty.
@@ Appendix F Empirical Estimation of Asymmetry Ratio from CAISO Market Data
F.1 Data Sources and Alignment (CAISO OASIS)
@@
F.2 Price-Spread Asymmetry (Robust Estimator)
@@
F.3 Calendar-2025 Empirical Estimates (Per Utility)
Table 9 reports empirical asymmetry ratios over calendar-2025 for each TAC area using representative settlement points (DLAPs or the NP15 trading hub). We report as the primary, bias-independent market-based estimator and include as an optional diagnostic when stable and available. Interpretation for risk-averse training. The implied can fall below in calendar-year averages; we do not interpret this as a recommendation to train below-median forecasts for reliability-critical operations. Instead, we treat as a market-grounded anchor and select an operationally conservative using an explicit reliability premium and bias-control constraints (Section 4.3).
@@ F.3 Calendar-2025 Empirical Estimates (Per Utility)
Notes: depends only on the DA–RT price spread and is therefore robust to forecast bias. conditions on the sign of the DA load forecast error and can be unstable or unavailable (reported as N/A) depending on data coverage and estimator settings. For risk-averse training, we treat market-derived asymmetry as an anchor and select an operationally conservative (Eq. 12) jointly with an explicit bias-control term (Section 4.3).
F.4 Event-Conditional Settlement Proxy (Optional)

### v6

@@ Appendix A Weather Covariates
Table IV lists the 8 meteorological covariates used in weather-integrated models. All weather data are sourced from NOAA Integrated Surface Database (ISD) stations within each utility service territory, aggregated to hourly resolution.
@@ Appendix B Hyperparameter Configuration
Table V provides complete hyperparameter specifications for all models.
| Parameter | S-Mamba | PowerMamba | Mamba-ProbTSF | PatchTST | iTransformer | LSTM |
| 128 | 128 | 128 | 256 | 512 | 256 | |
| State dim () | 16 | 16 | 16 | — | — | — |
| Conv kernel () | 4 | 4 | 4 | — | — | — |
| Patch length | — | — | — | 16 | — | — |
| Encoder layers | 2 | 2 | 2 | 3 | 3 | 2 |
| Attention heads | — | — | — | 4 | 8 | — |
| Dropout | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 | 0.1 |
| Bidirectional | Yes | Yes | Yes | — | — | Yes |
| Total parameters | 16.4M | 2.5M | 16.4M | 2.0M | 6.5M | 2.6M |
@@ Appendix C Multi-Seed Robustness
To assess robustness to random initialization, we trained PowerMamba + Weather with seeds . The 24-hour MAPE on CA ISO-TAC was (mean std), confirming that reported results are stable across initializations. These results confirm that the performance advantages of the evaluated architectures are robust to initialization variance.
@@ Appendix D Per-Utility Results
Table VI presents 24-hour MAPE for each TAC area, demonstrating consistent performance across diverse grid regions.
@@
Smaller utilities (SDGE-TAC, TIDC) exhibit higher MAPE due to increased load volatility relative to system size. Weather integration tends to improve accuracy, with the largest relative gains in smaller, more volatile territories; however, improvements are not uniform across all regions (Table VI).
@@ Appendix E Sensitivity to Weather Forecast Errors
Table VII demonstrates that the model retains its performance advantage even with moderate forecast errors. At (a typical error range for day-ahead weather forecasts), the performance degrades by less than 10%, remaining competitive with the non-weather-integrated baselines. This confirms that the benefits of weather integration largely persist under operational weather uncertainty.
@@ Appendix F Empirical Estimation of Asymmetry Ratio from CAISO Market Data
F-A Data Sources and Alignment (CAISO OASIS)
@@
F-B Price-Spread Asymmetry (Robust Estimator)
@@
F-C Calendar-2025 Empirical Estimates (Per Utility)
Table VIII reports empirical asymmetry ratios over calendar-2025 for each TAC area using representative settlement points (DLAPs or the NP15 trading hub). We report as the primary, bias-independent market-based estimator and include as an optional diagnostic when stable and available. Interpretation for risk-averse training. The implied can fall below in calendar-year averages; we do not interpret this as a recommendation to train below-median forecasts for reliability-critical operations. Instead, we treat as a market-grounded anchor and select an operationally conservative using an explicit reliability premium and bias-control constraints (Section IV-D).
@@ F-C Calendar-2025 Empirical Estimates (Per Utility)
Notes: depends only on the DA–RT price spread and is therefore robust to forecast bias. conditions on the sign of the DA load forecast error and can be unstable or unavailable (reported as N/A) depending on data coverage and estimator settings. For risk-averse training, we treat market-derived asymmetry as an anchor and select an operationally conservative (Eq. 12) jointly with an explicit bias-control term (Section IV-D).
F-D Event-Conditional Settlement Proxy (Optional)